OpenStudy (anonymous):

The radius of a cylinder is 15 cm, and its volume, found using 3.14 to approximate pi, is 19,782 cm3. What is the height of the cylinder? _____cm

OpenStudy (anonymous):

\(V=\pi r^2 h\)

OpenStudy (anonymous):

= pi 15^2 times 19,782? @surjithayer

OpenStudy (anonymous):

@jordanloveangel

OpenStudy (anonymous):

@mtwizzle420

OpenStudy (mtwizzle420):

Im no good a math.. :(

OpenStudy (anonymous):

Aww okay thanxx anyway

OpenStudy (anonymous):

@dtan5457

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

@MARC_

OpenStudy (anonymous):

@brookeutterback

OpenStudy (anonymous):

\[\pi *15^2*h=19,782,find~h\]